Find the value of the six trigonometric functions for an angle 0 in standard position if point P(-5,5sqrt(3))is on its terminal side.


I need help with this please! Thanks!

Answer by stanbon(75887)

Find the value of the six trigonometric functions for an angle 0 in standard position if point P(-5,5sqrt(3))is on its terminal side.
-------
Plot the point.
Notice it in QII.
-----------------------
You have x = -5 and y = 5sqrt(3)
So, r = sqrt[5^2 + (5sqrt(3))^2]
r = sqrt[25 + 75] = 10
---------------
sin = y/r = 5sqrt(3)/10 = (1/2)sqrt(3)
cos = x/r = -5/10 = -1/2
tan = y/x = 5sqrt(3)/-5 = -sqrt(3)
-----
csc = r/y = 2/sqrt(3)
sec = r/x = -2
cot = x/y = -1/sqrt(3)
==============================
cheers,
Stan H.
===========================